OpenStudy (anonymous):

The radius of a circle is 4 kilometers. What is the circle's circumference?

OpenStudy (anonymous):

the circumference of a circle is calculated by 2r, where r is the radius. what do you think the circumference is?

OpenStudy (anonymous):

i have no idea thats why im asking you guys

OpenStudy (anonymous):

8 maybe idk

OpenStudy (anonymous):

exactly ^.^ you got it

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

oh pellet, sorry haha its actually 2\[\pi \pi\]

OpenStudy (anonymous):

grr that didn't do it correctly. the formula is \[2pir\] so the answer is actually \[8\pi\] sorry about that!

OpenStudy (anonymous):

The diameter of a circle is 4 kilometers. What is the circle's circumference?

OpenStudy (anonymous):

the diameter is just 2r, so if the diameter is 4, divide it by 2 to get the radius once you have that, plug it into the circumference formula to get your answer \[2(2)\pi\] or you might notice that you're doing an extra step unnecessarily, and you can just see that the forumla is also \[c=D \pi\] where d is the diameter

OpenStudy (anonymous):

The diameter of a circle is 6 kilometers. What is the circle's circumference? Sorry im asking a lot if u have ever heard of IXL i hae to do it up till 90 or 80

OpenStudy (anonymous):

The radius of a circle is 3 centimeters. What is the circle's circumference? nevermind that one

OpenStudy (anonymous):

using the formula i just gave you, what do you think it is?

OpenStudy (anonymous):

so 2rPi, when r = 3?